# John Mendonça1,2, Isabel Trancoso1,2    Alon Lavie3,4

### v2

@@ D.2 Annotation Details
D.3 Additional analysis
Figure 8: Overall Score distribution across languages.
This behaviour is also evidenced when looking at the distribution of overall scores by language in Figure 8. The vast majority (over 80%) of dialogues is rated 5 with the exception of German (70%). We also note that no dialogue was rated with the worst score (1). This could be due to two distinct reasons: firstly, that LLM-chatbots are robust enough to recover from problematic conversations; secondly, we designed our LLM to act as a collaborative user, which incentives chatbots to recover.
Appendix E Benchmark Curation
E.1 Example Selection
Given our annotator budget, we were limited to annotating 100 dialogues per language. Given this relatively small sample size, we had to ensure our benchmark was composed of meaningful examples that represent the types of issues LLM-based chatbots incur. As such, we employ a sampling strategy that attempts to balance the distribution of quality in the benchmark.
We obtain this subset by randomly sampling one dialogue per user-chatbot that contains no issues and has a overall score of 5. For the remaining dialogues, we iterate over the types of issues until we obtain our desired benchmark size of 100 examples. For each iteration, we keep track of the distribution of overall score and chatbots and select a dialogue that increases the representation of the minority class. If the minority class is tied, we select one at random. If no examples that match the requirements are found, then we ignore that requirement in the selection  [^14] This was frequent when looking for dialogues from stronger chatbots with rarer issues..
@@ E.2.1 Overall Quality
We present Spearman correlations across languages for Overall Quality in Figure 9. We observe differences in performance across languages: correlations are highest in Portuguese and lowest in French. We also note that Deepspeek-V3 underperforms significantly in Spanish and French when compared to other models, while GPT-4.1 and O3 generally outperform other models.
@@ E.2.2 Issue Detection
For completeness, we report F1+ scores (Figure 11) for each individual language. Confounding factors – such as variation in chatbot performance across languages and annotator-specific interpretations or nuances – make a language-wise analysis less straightforward for determining cross-lingual performance of LLMs as evaluators.
@@ E.3 Additional Examples
We present additional examples from our curated meta-evaluation benchmark in Figures 12, 13, 14 and 15.

### v3

@@ D.1 Annotation Details
D.2 Additional analysis
This behaviour is also evidenced when looking at the distribution of overall scores by language in Figure 9. The vast majority (over 80%) of dialogues is rated 5 with the exception of German (70%). We also note that no dialogue was rated with the worst score (1). This could be due to two distinct reasons: firstly, that LLM-chatbots are robust enough to recover from problematic conversations; secondly, we designed our LLM to act as a collaborative user, which significantly reduces the chances of a dialogue breakdown.
Language !Int !Safe !Emp !Com Rep !Coh !Rel !Fac Overall
EN 1.000 .6414 .4792 .5581 .2803 .5885 .4501 .7361 .4707
ZH .3258 .8236 .5415 .5811 .3981 .4236 .3220 .8347 .2622
FR .3716 .5811 .3877 .6581 .3877 .4212 .3886 .7967 .3508
PT .3337 .5522 .3926 .4243 .2908 .2419 .2562 .8415 .4182
Table 3: Inter-Annotator Agreement (IAA) for all dimensions across languages with 2 annotators. A formal definition of these dimensions is available in Table 1.
Appendix E Benchmark Curation
Figure 9: Overall Score distribution across languages.
Figure 10: Pearson correlations across languages (all significant with $p<0.01$). Combined denotes the benchmark composed by all individual languages.
E.1 Human Annotations
We present Inter-Annotator Agreement (IAA) agreement scores in Table 3, calculated with using Krippendorff’s $\alpha$ [^13] We omit bootstrap intervals due to instability caused by the low number of positive examples in an overall small sample size (100), which leads to unreliable or undefined CI.. Issues such as non-factual (!Fac) and unsafe (!Safe) achieve very high correlations. This is due to the relatively objective nature of this issue [^14] Furthermore, annotators were allowed to consult external references to check for ”factuality”.. Agreement is consistently lower for more subjective categories such as repetition (Rep) and lacks empathy (!Emp), highlighting the challenges of reliably annotating these issues.
@@ E.2.1 Overall Quality
We present per-language Pearson correlations for Overall Quality in Figure 10 and numeric correlations results in Table 4.
@@ E.2.2 Issue Detection
Model Pearson Spearman
GPT-4.1 .4767 .4525
Gemini-2.0-flash .3707 .3558
Deepseek-V3 .3206 .2817
Deepseek-R1 .4472 .4044
Gemini-2.5-flash .4462 .4200
o3 .4914 .4750
Table 4: Numeric overall quality correlation results on the full benchmark. All correlation results are significant with $p<0.01$. Bold denotes best model.
@@
For completeness, we report F1+ scores (Figure 13) for each individual language.
@@ E.3 Additional Examples
We present additional examples from our curated meta-evaluation benchmark in Figures 14, 15, 16 and 17.